Present value. The State of Confusion wants to change the current retirement policy for state employees. To do so, however, the state must pay the current pension fund members the present value of their promised future payments. There are 260,000 current employees in the state pension fund. The average employee is 20 years away from retirement, and the average promised future retirement benefit is $350,000 per employee. If the state has a discount rate of 5% on all its funds, how much money will the state have to pay to the employees before it can start a new pension plan?
Answer #1

FV of the retirement payment after 20 years = 350,000 per employee
PV of The benefit = 350,000/(1+5%)

Total Amount to be paid to all its employees = 260,000 * 350,000/1.05 = 86,666,666.66

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well